Sutton Products is a

priceminus−setter

that uses the

costminus−plus

pricing approach. The products are specialty vacuum tubes used in sound equipment. The CEO is certain that the company can produce and sell

300 comma 000300,000

units per​ year, due to the high demand for the product. Variable costs are

$ 2.30$2.30

per unit. Total fixed costs are

$ 970 comma 000$970,000

per year. The CEO will receive stock options if

$ 400 comma 000$400,000

of operating income for the year is reported. What sales price would allow the CEO to achieve the target if the

costminus−plus

pricing method is​ used? (Round your answer to the nearest​ cent.)

A.

$ 4.57$4.57

per unit

B.

$ 2.30$2.30

per unit

C.

$ 4.20$4.20

per unit

D.

$ 6.87$6.87

per unit
Answer #1

Total cost of the product = Variable cost + Fixed cost

= (300000*2.30+970000) = $1660000

Sales price = (Total cost of the product + Desired operating profit) / No. of units

= (1660000+400000)/300000

= $6.87 per unit

Option D. is correct answer.